# The Discrepancy of Shortest Paths

### v2

@@ A.1 Discrepancy on graphs
In graph theory, the discrepancy of a graph introduced by Erdős [Erd63] is defined as follows:
@@
where $e(G[U])$ is the number of edges of the induced subgraph $G[U]$ on vertices $U$ and $p=|E|/{n\choose 2}$ is the density of edges. If we consider a complete graph and randomly color each edge with probability $p$, the above definition of discrepancy quantifies the deviation of induced subgraphs of $G$ from their expected size. Erdős and Spencer [ES72] showed that the graph discrepancy is $\Theta(n^{3/2})$ when $p=1/2$. This definition and related definitions (e.g., positive discrepancy, dropping the absolute operator) have applications to quasi-randomness [CGW89], graph cuts and edge expansion [Alo97, BS06, RST23]. There is also study of multicolor discrepancy [GKM22, FHLT21] that we skip here.
Of particular relevance to our work, Balogh et al. [BCJP20] studied edge discrepancy (as defined in this paper) of (spanning) trees, paths and Hamilton cycles of a graph $G$. In particular, they showed that, for any labeling of edges of $G$, there is a path with discrepancy $\Omega(n)$, even when the graph is a grid. Prior to this, either probabilistic construction exhibiting such a lower bound was known [Bec83, DP15, Let16] or an explicit construction of linear size non-planar graphs [AC88], resolving a question by Erdős [Erd63]. The construction for the planar graph in [BCJP20] can be extended to coloring of vertices such that there is a path with vertex discrepancy $\Omega(n)$ when there are exponentially many paths and $\Omega(\sqrt{n})$ when there are polynomially many paths. (see Appendix C for details).
Discrepany of paths in directed graphs has also been studied. Reimer [Rei02] showed that, if a directed graph has discrepancy $\Omega(n)$, then the graph must have $\Omega(n^{2})$ edges. In the case when we do not allow antiparallel edges, Ben-Eliezer et al. [BEKS12] showed that there is a directed graph with $\Theta(n^{2}\log^{2}(n))$ edges such that any mapping $\chi:E\to\{-1,1\}$ will either have a path of length $\Omega(n)$ and all edges mapped to $-1$ or a path of length $\Omega(n\log(n))$ with all edges mapped to $+1$.
@@ A.2 Connection with curve discrepancy
A classical topic in computational geometry is to study upper and lower bounds of the discrepancy/hereditary discrepancy of the incidence matrix of geometric objects and a set of points. For example, for a set of $n$ points and $n$ halfplanes in $\mathbb{R}^{2}$, the $n$ by $n$ incidence matrix (with rows corresponding to halfplanes and columns corresponding to points) has discrepancy of $\Omega(n^{1/4})$. For $n$ points and $n$ lines in the plane, the discrepancy of the incidence matrix is $\Omega(n^{1/6})$ [CL00]. In general the discrepancy of such incidence matrix is related to the ‘complexity’ of the geometric shapes. In our setting of a graph, the set of all pairs shortest paths defines a set system on the vertices. When the graph is planar, the shortest paths are essentially simple curves in the plane.
We would like to compare our results with discrepancy of curves and points in the geometric setting. Using the classification in [PS98], a family of simple curves have $k$ degree of freedom and multiplicity type $s$ if for any $k$ points there are at most $s$ curves passing through all of them, and any pair of curves intersect in at most $s$ points. Lines in the plane have degree of $2$ and multiplicity of $1$. A set of curves with degree of $2$ and multiplicity of $1$ is called pseudolines – two pseudolines have at most one intersection. For $n$ points and $n$ lines, the discrepancy is upper bounded by $O(n^{1/6}(\log n)^{2/3})$ [Cha00] which is nearly tight by a polylogarithmic factor. The proof uses the standard partial coloring argument with the Szeremédi-Trotter bound on point-line incidence – for any $n$ points and $m$ lines there are at most $O(m^{2/3}n^{2/3}+m+n)$ point-line incidences [ST83]. The Szeremédi-Trotter bound can be extended to a set of pseudolines [PS98, Szé97]. Therefore the same proof and upper bound hold for the discrepancy of pseudolines.
For a consistent set of shortest paths, two shortest paths will only intersect at a contiguous segment, which may have multiple vertices/points. Thus using the curve classification criterion, a consistent family of shortest paths in the plane has degree of $2$ but multiplicity $s$ that is possibly higher than a constant. In fact, our discrepancy lower bound construction in the planar graph setting uses a design with $s$ possibly as high as $n^{1/2}$. This is the major difference of shortest paths in a planar graph with pseudolines, which allows the discrepancy of shortest paths to go beyond the pseudoline upper bound of $\widetilde{O}(n^{1/6})$ [^6] Using the incidence upper bound for $k=2$ and $s=n^{1/2}$ from [Szé97] and partial coloring, one can obtain a discrepancy upper bound of $\widetilde{O}(n^{1/3})$ for our path construction. In contrast, we obtained nearly tight bound of $\widetilde{\Theta}(n^{1/4})$..
@@ Appendix B Technical Preliminaries
We first observe that, using a random coloring $\chi$, we can guarantee that for all paths $\pi\in\Pi$ [Cha00]:
@@
We have defined vertex and edge (hereditary) discrepancy; one may wonder if there is an underlying relationship between vertex and edge (hereditary) discrepancy since they share the same bounds in most scenarios as presented in Table 1. The following observation shows that vertex discrepancy directly implies bounds on edge discrepancy.
@@
For the lower bound on hereditrary discrepancy, one general result is the trace bound [CL00].
Lemma B.5 (Trace Bound [CL00]).
@@
Recently, this trace bound has been improved in a series of works culminating in the following bound in Larsen [Lar19] which we also use in our work:
Lemma B.6 (Trace Bound [Lar19]).
@@
Appendix C Discrepancy Bounds for Paths Without Consistency
@@
The first claim follows from Proposition 1.6 in [BCJP20], which says that the edge discrepancy of paths on a $k\times\ell$ grid graph is at least $\Omega(k\ell)$. To prove vertex discrepancy, we make two additional remarks about this construction. First, for our purpose, it is sufficient to consider only an $n\times 2$ grid graph $G$. The set of paths in the construction of [BCJP20] consists of all paths that start from the top left corner and bottom left corner going to the right and possibly taking a subset of the vertical edges in the grid graph. The number of paths is $O(2^{n})$. Second, for vertex discrepancy, we define a companion graph $G^{\prime}$. Specifically, for each grid edge $e$ in $G$, we place a vertex $v$ of $G^{\prime}$ on $e$. We connect two vertices in $G^{\prime}$ if and only if the corresponding edges in $G$ share a common vertex. The graph $G^{\prime}$ is still planar. In additional, a path $P$ in $G$ maps to a corresponding path $P^{\prime}$ in $G^{\prime}$ where vertices on $P^{\prime}$ follow the same order of the corresponding edges on $P$. See Figure 5 for an example. Therefore the edge discrepancy in $G$ and the vertex discrepancy of $G^{\prime}$ are the same.
@@
For the second claim, we take an $n\times n$ Hadamard matrix $H$ with $n$ as power of $2$. The elements in $H$ are $+1$ or $-1$. Each row has $n/2$ elements of $+1$ and $n/2$ elements of $-1$ and the rows are pairwise orthogonal. It is known [Cha00] that the matrix $A=\{a_{ij}\}=\frac{1}{2}(H+J)$ with $J$ as an $n\times n$ matrix of all $1$ has discrepancy at least $\Omega(\sqrt{n})$.
@@
Appendix D Relation Between Discrepancy and Hereditary Discrepancy in General Graphs
In Theorem 5, we showed a construction of a path weighted graph $G$ whose system of unique shortest paths $\Pi$ satisfies $\operatorname{herdisc}_{v}(\Pi)\geq\Omega(n^{1/4}/\sqrt{\log(n)})$. Here, we will observe that this result extends to discrepancy:
@@ Appendix D Relation Between Discrepancy and Hereditary Discrepancy in General Graphs
Let $G$ be the graph from Theorem 5, and let $\Pi$ be its system of unique shortest paths. By definition of hereditary discrepancy, there exists an induced path subsystem $\Pi^{\prime}\subseteq\Pi$ with
@@
It thus suffices to argue that there is a graph $G^{\prime}$ on $n^{\prime}\leq n$ nodes in which all paths in $\Pi^{\prime}$ are unique shortest paths. Indeed, this is well known, and is shown e.g. in [Bod19] (c.f. Lemma 2.4.4 and 2.4.11). To sketch the proof: suppose that a node $v$ is deleted from the initial system $\Pi$. Consider each path $\pi\in\Pi$ that contains $v$ as an internal node, i.e., it has the form
@@
Appendix E Application in Matrix Analysis
@@
One can write eq 2 in a form of a semi-definite program (see Lee et al. [LSŠ08]) and also show that the Slater point exists. In particular, the primal and dual program coincides. An interesting question in matrix analysis is to estimate the factorization norm of different class of matrices. In a series of work, various authors have computed tight bounds on the factorization norm of certain class of matrices:
  • If $A$ is a unitary matrix, then $\gamma_{2}(A)=1$ [Sch11].
  • If $A\in\mathbb{R}^{n\times n}$ is a positive semidefinite matrix with entries $A_{ij}$ [Sch11], then
@@
  • Mathias [Mat93] $A$ satisfies the property that $\sqrt{A^{\top}A}\bullet\mathbb{I}=\sqrt{AA^{\top}}\bullet\mathbb{I}={\operatorname{tr}(A)\over n}\mathbb{I}$, then
@@
    In particular, if $A\in\{0,1\}^{n}$ is a lower-triangular one $1$ matrix, then $\gamma_{2}(A)=\Theta(\log n)$ [FHU23, HUU23, KP70].
  • If $A\in\mathbb{R}^{n\times n}$ is a lower-triangular Toeplitz matrix with entries decreasing either polynomially or exponentially, then $\gamma_{2}(A)=\Theta(1)$ [HUU24].
@@
Lemma E.1 (Matoušek et al. [MNT20]).
@@
Combining Lemma E.1 with our results, we have the following corollary:
Corollary E.2.
@@
Appendix F Differentially Private All Sets Range Queries
Here we introduce the DP-ASRQ problem and show its connection to the DP-APSD problem.
Given an undirected graph $G$, the problem of All Sets Range Queries (ASRQ) considers each edge associated with a certain attribute, and the range is the set of edges along a shortest path. Two type of queries are considered here: the bottleneck query returns the largest/smallest attribute in a range; while the counting query returns the summation of all attributes.
At a schematic level, the ASRQ problem with counting queries is very similar to the APSD problem: the graph topology is public and the edge attributes are considered private. Only that the shortest path structure is dictated by the edge weights to be protected in the APSD problem, however, irrelevant to the edge attributes in the ASRQ problem. This subtle difference has consequential caveat in the algorithm design: the graph topology can be used, for example, to construct an exact hopset first then apply perturbations to the edge attributes in the ASRQ problem; nevertheless, approach of this kind violates protecting the sensitive information in the APSD problem, since adversarial inference can be made on edge weights when the graph topology information is used. This observation also notes that the APSD problem is strictly harder than the ASRQ problem: recall that the best additive error upper bound of the APSD problem is still $\widetilde{O}(n^{1/2})$, while the other has $\widetilde{O}(n^{1/4})$ [DGUW23]. Therefore, plugging in our new hereditary lower bound essentially closes the gap for the ASRQ problem.
Corollary F.1 (Formal version of Theorem 4).
Given an $n$-node undirected graph, for any $\beta\in(0,1)$ and any $\varepsilon,\delta>0$, no $(\varepsilon,\delta)$-DP algorithm for ASRQ has additive error of $o(n^{1/4})$ with probability $1-\beta$.
Definition 7 (Neighboring weights [Sea16]).
For a graph $G=(V,E)$, let $w,w^{\prime}:E\rightarrow\mathbb{R}^{\geq 0}$ be two weight functions that map any $e\in E$ to a non-negative real number, we say $w,w^{\prime}$ are neighboring, denoted as $w\sim w^{\prime}$ if $\sum_{e\in E}|w(e)-w^{\prime}(e)|\leq 1.$
Definition 8 (Differentially Private Range Queries).
Let $(\mathcal{R}=(X,\mathcal{S}),f)$ be a system of range queries and $w,w^{\prime}:X\rightarrow\mathbb{R}^{\geq 0}$ be neighboring attribute functions. Furthermore, let $\mathcal{A}$ be an algorithm that takes $(\mathcal{R},f,w)$ as input. Then $\mathcal{A}$ is $(\varepsilon,\delta)$-differentially private on $G$ if, for all pairs of neighboring attribute functions $w,w^{\prime}$ and all sets of possible outputs $\mathcal{C}$, we have $\Pr[\mathcal{A}(\mathcal{R},f,w)\in\mathcal{C}]\leq e^{\varepsilon}\cdot\Pr[\mathcal{A}(\mathcal{R},f,w^{\prime})\in\mathcal{C}]+\delta.$ If $\delta=0$, we say $\mathcal{A}$ is $\varepsilon$-differentially private on $G$.
To complete this section, we give the formal definition of the ASRQ problem above. The definition of neighboring attributes follows Definition 7. The lower bound proof of Corollary F.1 simply imitates the APSD problem, because the reduction from the linear query problem still holds despite the difference between two problems. The proof is omitted to avoid redundancy.

### v3

@@ A.1 Discrepancy on graphs
In graph theory, the discrepancy of a graph introduced by Erdős [erdHos1963combinatorial] is defined as follows:
@@
where $e(G[U])$ is the number of edges of the induced subgraph $G[U]$ on vertices $U$ and $p=|E|/{n\choose 2}$ is the density of edges. If we consider a complete graph and randomly color each edge with probability $p$, the above definition of discrepancy quantifies the deviation of induced subgraphs of $G$ from their expected size. Erdős and Spencer [erdiis1972imbalances] showed that the graph discrepancy is $\Theta(n^{3/2})$ when $p=1/2$. This definition and related definitions (e.g., positive discrepancy, dropping the absolute operator) have applications to quasi-randomness [chung1989quasi], graph cuts and edge expansion [alon1997edge, bollobas2006discrepancy, raty2023positive]. There is also study of multicolor discrepancy [gishboliner2022discrepancies, freschi2021note] that we skip here.
Of particular relevance to our work, Balogh et al. [balogh2020discrepancies] studied edge discrepancy (as defined in this paper) of (spanning) trees, paths and Hamilton cycles of a graph $G$. In particular, they showed that, for any labeling of edges of $G$, there is a path with discrepancy $\Omega(n)$, even when the graph is a grid. Prior to this, either probabilistic construction exhibiting such a lower bound was known [beck1983size, dudek2015alternative, letzter2016path] or an explicit construction of linear size non-planar graphs [alon1988explicit], resolving a question by Erdős [erdHos1963combinatorial]. The construction for the planar graph in [balogh2020discrepancies] can be extended to coloring of vertices such that there is a path with vertex discrepancy $\Omega(n)$ when there are exponentially many paths and $\Omega(\sqrt{n})$ when there are polynomially many paths. (see Appendix D for details).
Discrepany of paths in directed graphs has also been studied. Reimer [reimer2002ramsey] showed that, if a directed graph has discrepancy $\Omega(n)$, then the graph must have $\Omega(n^{2})$ edges. In the case when we do not allow antiparallel edges, Ben-Eliezer et al. [ben2012size] showed that there is a directed graph with $\Theta(n^{2}\log^{2}(n))$ edges such that any mapping $\chi:E\to\{-1,1\}$ will either have a path of length $\Omega(n)$ and all edges mapped to $-1$ or a path of length $\Omega(n\log(n))$ with all edges mapped to $+1$.
@@ A.2 Connection with curve discrepancy
A classical topic in computational geometry is to study upper and lower bounds of the discrepancy/hereditary discrepancy of the incidence matrix of geometric objects and a set of points. For example, for a set of $n$ points and $n$ halfplanes in $\mathbb{R}^{2}$, the $n$ by $n$ incidence matrix (with rows corresponding to halfplanes and columns corresponding to points) has discrepancy of $\Omega(n^{1/4})$. For $n$ points and $n$ lines in the plane, the discrepancy of the incidence matrix is $\Omega(n^{1/6})$ [Chazelle2000-ld]. In general the discrepancy of such incidence matrix is related to the ‘complexity’ of the geometric shapes. In our setting of a graph, the set of all pairs shortest paths defines a set system on the vertices. When the graph is planar, the shortest paths are essentially simple curves in the plane.
We would like to compare our results with discrepancy of curves and points in the geometric setting. Using the classification in [Pach1998-ib], a family of simple curves have $k$ degree of freedom and multiplicity type $s$ if for any $k$ points there are at most $s$ curves passing through all of them, and any pair of curves intersect in at most $s$ points. Lines in the plane have degree of $2$ and multiplicity of $1$. A set of curves with degree of $2$ and multiplicity of $1$ is called pseudolines – two pseudolines have at most one intersection. For $n$ points and $n$ lines, the discrepancy is upper bounded by $O(n^{1/6}(\log n)^{2/3})$ [Chazelle2000-rd] which is nearly tight by a polylogarithmic factor. The proof uses the standard partial coloring argument with the Szeremédi-Trotter bound on point-line incidence – for any $n$ points and $m$ lines there are at most $O(m^{2/3}n^{2/3}+m+n)$ point-line incidences [Szemeredi1983-tv]. The Szeremédi-Trotter bound can be extended to a set of pseudolines [Pach1998-ib, Szekely1997-yp]. Therefore the same proof and upper bound hold for the discrepancy of pseudolines.
For a consistent set of shortest paths, two shortest paths will only intersect at a contiguous segment, which may have multiple vertices/points. Thus using the curve classification criterion, a consistent family of shortest paths in the plane has degree of $2$ but multiplicity $s$ that is possibly higher than a constant. In fact, our discrepancy lower bound construction in the planar graph setting uses a design with $s$ possibly as high as $n^{1/2}$. This is the major difference of shortest paths in a planar graph with pseudolines, which allows the discrepancy of shortest paths to go beyond the pseudoline upper bound of $\widetilde{O}(n^{1/6})$ [^1] Using the incidence upper bound for $k=2$ and $s=n^{1/2}$ from [Szekely1997-yp] and partial coloring, one can obtain a discrepancy upper bound of $\widetilde{O}(n^{1/3})$ for our path construction. In contrast, we obtained nearly tight bound of $\widetilde{\Theta}(n^{1/4})$..
@@ Appendix B Technical Preliminaries
We first observe that, using a random coloring $\chi$, we can guarantee that for all paths $\pi\in\Pi$ [Chazelle2000-rd]:
@@
We have defined vertex and edge (hereditary) discrepancy; one may wonder if there is an underlying relationship between vertex and edge (hereditary) discrepancy since they share the same bounds in most scenarios as presented in LABEL:tab:res-summary. The following observation shows that vertex discrepancy directly implies bounds on edge discrepancy.
@@
For the lower bound on hereditrary discrepancy, one general result is the trace bound [Chazelle2000-ld].
Lemma B.5 (Trace Bound [Chazelle2000-ld]).
@@
Recently, this trace bound has been improved in a series of works culminating in the following bound in Larsen [larsen2017constructive] which we also use in our work:
Lemma B.6 (Trace Bound [larsen2017constructive]).
@@
Appendix C Missing Proofs from LABEL:subsec:edge-disc-exp
Proof of LABEL:prop:disc_E_1.
The proposition follow from an argument similar to Proposition LABEL:prop:disc_bound. For each path $\pi^{\prime}\in\Pi^{\prime}$, let $X_{\pi^{\prime}}$ be the random variable defined as
We can restate the discrepancy of $E_{1}$ as
By Proposition LABEL:prop:edge_small_disc, if $|E(\pi)\cap E(\pi^{\prime})|\equiv 0\mod 2$, then $X_{\pi^{\prime}}=0$, so without any loss of generality, we may assume that $|E(\pi)\cap E(\pi^{\prime})|$ is odd for all $\pi^{\prime}\in\Pi^{\prime}$. In this case,
implying that $\mathbb{E}[\sum_{\pi^{\prime}\in\Pi^{\prime}}X_{\pi^{\prime}}]=0$. Then by Proposition LABEL:prop:edge_path_cover and the Chernoff bound, it follows that for any constant $c\geq 1$,
Proof of LABEL:prop:disc_E_2.
The proposition follow from an argument similar to Proposition LABEL:prop:disc_bound. Note that by the Edge Covering Property of the path cover (Proposition LABEL:prop:edge_path_cover),
completing the proof. ∎
Proof of LABEL:prop:disc_E_3.
Let
denote the number of paths in our path cover that intersect $\pi$. We define a function $f:\mathbb{Z}_{\geq 0}\mapsto\mathbb{Z}_{\geq 0}$ such that $f(\phi)$ equals the largest possible value of $|E_{3}|$ when $\phi=k$. Note that $f$ is well-defined since $0\leq|E_{3}|\leq|E|$. We will prove that $f(\phi)\leq 4\phi,$ by recursively decomposing path $\pi$. When $\phi=1$, there is only one path $\pi^{\prime}\in\Pi^{\prime}$ that intersects $\pi$. Then the only edges in $E_{3}$ are of the form
By the Consistency Property of Proposition LABEL:prop:edge_path_cover, path $\pi^{\prime}$ can intersect $\pi$ and then split apart at most once. Then
When $\phi>1$, we will split our analysis into the two cases:
    $\pi=\pi[s,v_{1}]\circ\pi[x,y]\circ\pi[v_{3},t],$
    where $\circ$ denotes the concatenation operation. Let
    We claim that $\phi_{1}<\phi$, $\phi_{2}<\phi$, and $\phi_{1}+\phi_{2}=\phi$. We will use these facts to establish a recurrence relation for $f$. By our assumption that $\pi[v_{1},v_{3}]\cap\pi_{1}^{\prime}=\{v_{1},v_{3}\}$, it follows that $\pi[x,y]\cap\pi_{1}^{\prime}=\emptyset$, and so $\phi_{1}<\phi$. Likewise, by the No Repeats Property of Proposition LABEL:prop:edge_path_cover,
    $(\pi[s,v_{1}]\circ\pi[v_{3},t])\cap\pi_{2}^{\prime}=\emptyset,$
    so $\phi_{2}<\phi$. Finally, observe that more generally, if there exists a path $\pi^{\prime}\in\Pi^{\prime}$ such that $\pi^{\prime}\cap\pi[x,y]\neq\emptyset$ and $\pi^{\prime}\cap(\pi[s,v_{1}]\circ\pi[v_{3},t])\neq\emptyset$, then the No Repeats Property of Proposition LABEL:prop:edge_path_cover is violated. We conclude that $\phi_{1}+\phi_{2}=\phi$.
    Now $|E_{3}|$ can be upper bounded by the following inequality:
    $f(\phi)\leq f(\phi_{1})+f(\phi_{2})+2=f(i)+f(\phi-i)+2,$
    where $0<i<\phi$.
    $\pi[s,t]=\pi[s,x]\circ\pi[v_{1},v_{2}]\circ\pi[y,t].$
    Let
    we immediately obtain the recurrence
    $f(\phi)\leq f(\phi_{1})+f(\phi_{2})+2\leq f(1)+f(\phi-1)+2.$
Applying this recurrence $\leq\phi$ times, we find that
$f(\phi)\leq\phi\cdot f(1)+2\phi\leq 4\phi.$
$|E_{3}|\leq f(k)\leq f(n^{1/2})=O(n^{1/2}).$
Since the edges in $E_{3}\subseteq E\setminus E^{\prime}$ are labeled independently at random, we may apply a Chernoff bound as in Proposition LABEL:prop:edge_disc_e2 to argue that $\chi(E_{3})=O(n^{1/4}\log^{1/2}n)$ with high probability. ∎
Appendix D Discrepancy Bounds for Paths Without Consistency
@@
The first claim follows from Proposition 1.6 in [balogh2020discrepancies], which says that the edge discrepancy of paths on a $k\times\ell$ grid graph is at least $\Omega(k\ell)$. To prove vertex discrepancy, we make two additional remarks about this construction. First, for our purpose, it is sufficient to consider only an $n\times 2$ grid graph $G$. The set of paths in the construction of [balogh2020discrepancies] consists of all paths that start from the top left corner and bottom left corner going to the right and possibly taking a subset of the vertical edges in the grid graph. The number of paths is $O(2^{n})$. Second, for vertex discrepancy, we define a companion graph $G^{\prime}$. Specifically, for each grid edge $e$ in $G$, we place a vertex $v$ of $G^{\prime}$ on $e$. We connect two vertices in $G^{\prime}$ if and only if the corresponding edges in $G$ share a common vertex. The graph $G^{\prime}$ is still planar. In additional, a path $P$ in $G$ maps to a corresponding path $P^{\prime}$ in $G^{\prime}$ where vertices on $P^{\prime}$ follow the same order of the corresponding edges on $P$. See Figure 1 for an example. Therefore the edge discrepancy in $G$ and the vertex discrepancy of $G^{\prime}$ are the same.
@@
For the second claim, we take an $n\times n$ Hadamard matrix $H$ with $n$ as power of $2$. The elements in $H$ are $+1$ or $-1$. Each row has $n/2$ elements of $+1$ and $n/2$ elements of $-1$ and the rows are pairwise orthogonal. It is known [Chazelle2000-rd] that the matrix $A=\{a_{ij}\}=\frac{1}{2}(H+J)$ with $J$ as an $n\times n$ matrix of all $1$ has discrepancy at least $\Omega(\sqrt{n})$.
@@
Appendix E Relation Between Discrepancy and Hereditary Discrepancy in General Graphs
In Theorem LABEL:thm:undi-hdisc-lb, we showed a construction of a path weighted graph $G$ whose system of unique shortest paths $\Pi$ satisfies $\operatorname{herdisc}_{v}(\Pi)\geq\Omega(n^{1/4}/\sqrt{\log(n)})$. Here, we will observe that this result extends to discrepancy:
@@ Appendix E Relation Between Discrepancy and Hereditary Discrepancy in General Graphs
Let $G$ be the graph from Theorem LABEL:thm:undi-hdisc-lb, and let $\Pi$ be its system of unique shortest paths. By definition of hereditary discrepancy, there exists an induced path subsystem $\Pi^{\prime}\subseteq\Pi$ with
@@
It thus suffices to argue that there is a graph $G^{\prime}$ on $n^{\prime}\leq n$ nodes in which all paths in $\Pi^{\prime}$ are unique shortest paths. Indeed, this is well known, and is shown e.g. in [Bodwin19] (c.f. Lemma 2.4.4 and 2.4.11). To sketch the proof: suppose that a node $v$ is deleted from the initial system $\Pi$. Consider each path $\pi\in\Pi$ that contains $v$ as an internal node, i.e., it has the form
@@
Appendix F Application in Matrix Analysis
@@
One can write eq 1 in a form of a semi-definite program (see Lee et al. [lee2008direct]) and also show that the Slater point exists. In particular, the primal and dual program coincides. An interesting question in matrix analysis is to estimate the factorization norm of different class of matrices. In a series of work, various authors have computed tight bounds on the factorization norm of certain class of matrices:
  • If $A$ is a unitary matrix, then $\gamma_{2}(A)=1$ [schur1911remarks].
  • If $A\in\mathbb{R}^{n\times n}$ is a positive semidefinite matrix with entries $A_{ij}$ [schur1911remarks], then
@@
  • Mathias [mathias1993hadamard] $A$ satisfies the property that $\sqrt{A^{\top}A}\bullet\mathbb{I}=\sqrt{AA^{\top}}\bullet\mathbb{I}={\operatorname{tr}(A)\over n}\mathbb{I}$, then
@@
    In particular, if $A\in\{0,1\}^{n}$ is a lower-triangular one $1$ matrix, then $\gamma_{2}(A)=\Theta(\log n)$ [fichtenberger2023constant, henzinger2023almost, kwapien1970main].
  • If $A\in\mathbb{R}^{n\times n}$ is a lower-triangular Toeplitz matrix with entries decreasing either polynomially or exponentially, then $\gamma_{2}(A)=\Theta(1)$ [henzinger2023unifying].
@@
Lemma F.1 (Matoušek et al. [matouvsek2020factorization]).
@@
Combining Lemma F.1 with our results, we have the following corollary:
Corollary F.2.